The vapor pressure of CCl3F at 300 K is 856 torr.

If 11.0 g of CCl3F is enclosed in a 1.2 L container, calculate the number of moles of CCl3F in the gas phase.

Express your answer using two significant figures.
Answer #1

Assuming all CCl3F (300 K,856 torr) in vapor to a ideal gas. So by applying PV= nRT

V =1.2L(Volume of liquid will be negligible as compared to vapor)

P = (856/760)atm

n = PV/RT = 856*1.2/(0.0821*300*760) = 0.054875 ~0.055

So the number of moles of CCl3F in the gas phase in significant figures = 0.055